radius tires tra km before wearing them out. How manyr do the tires make, any backing up and any change in radus due to revolutions
Answer #1

radius of a tyre, R= 0.260 m

Distance covered in one revolution/circumference of a tyre, C= 2*pi*R

= 2*3.14*0.260

= 1.6328 m

Number of revolution ,n = total distance (in meters)/ C

= 70000000/1.6328

= 42871141 (approx)